# Visual and Memory Dual Adapter for Multi-Modal Object Tracking

### v1

@@ IV-C Comparison with the State-of-the-Art Methods
To evaluate the impact of different components on the tunable parameter and model complexity, we conduct a comparison of tunable parameters on the LasHeR dataset. As shown in Table VI, the first row represents the baseline model, where all components except the patch-embedding layer are frozen, resulting in a total of 93.4M parameters, with 1.9M being tunable. The second row shows the result of making all components of the baseline tunable, which improves the tracking performance by 4.3% in precision rate and 3.7% in success rates, however, the number of tunable parameters rises to 93.4M. In the third row, the proposed visual adapter is added to the baseline and made learnable, adding only 1.9M tunable parameters, while improving the tracking performance by 5.9% in precision rate and 4.7% in success rate, compared to the baseline. The last row shows the final version of the proposed method, which has 98.9M model parameters and 7.3M of them are tunable.
IV-H Comparison of Visual Adapters.
@@
We conduct comparison experiments between our proposed visual adapter and the widely adopted Fovea fusion adapter [24, 47, 55], fine-tuning both under identical settings. The “–” row in Table V represents a strong baseline without any dedicated adapter. Introducing the Fovea fusion adapter improves tracking performance on the LasHeR dataset, with precision and success rates increasing from 65.9% and 51.8% to 69.0% and 54.4%, respectively. Similar trends are observed on DepthTrack and VisEvent, where Fovea yields modest but consistent gains, indicating that spatial re-weighting of features contributes positively to performance. Replacing Fovea with our frequency-aware visual adapter leads to further improvements, achieving 71.8% precision and 56.5% success on LasHeR. This represents absolute gains of +5.9% (PR) and +4.7% (SR) over the baseline, and +2.8% / +2.1% over Fovea. A similar trend is seen on DepthTrack, where the F-score improves from 59.8% (no adapter) and 60.4% (Fovea) to 61.3%. On VisEvent, precision rises from 78.4% (baseline) and 79.0% (Fovea) to 79.5%. Although our module introduces 3.8 M tunable parameters in total, this is only 1.7 M more than the Fovea variant and remains lightweight compared with full fine-tuning of the backbone. The extra capacity is used to model frequency, domain interactions across RGB and auxiliary modalities, which explains why the gain is most pronounced on LasHeR, where motion blur and illumination variation make high-frequency cues especially informative, while still yielding steady improvements on the other two datasets. In sum, Table V confirms that our adapter offers the best accuracy–efficiency trade-off among the tested designs, outperforming both the plain baseline and the strong Fovea fusion across all metrics without incurring a prohibitive parameter cost.
@@
IV-I Qualitative Comparison and Feature Analysis.
@@
2) Feature Representation Analysis: To provide a more comprehensive visual analysis, we further inspect the feature visualization produced by both the visual and memory adapters.
For the visual adapter, Figure 8 arranges three columns: the left column shows the raw search region, the middle column depicts the feature map generated by the baseline without our adapter, and the right column presents the features after using the adapter. Once the adapter is enabled, the target contour becomes crisp and the clutter in the background is markedly suppressed, confirming that frequency-aware recalibration strengthens discriminative cues while filtering out irrelevant noise.
The memory adapter cannot be plotted directly because it stores one-dimensional vectors, so we instead compare the target‐region feature maps before and after the memory adapter in Figure 9. After memory aggregation, the foreground activations are noticeably sharper and the separation from the background is clearer, illustrating that temporal information preserved in memory further refines spatial representations and stabilizes tracking under appearance changes.
V Conclusion
In this paper, we proposed a novel prompt-learning-based tracker that combines visual and memory adapters for boosting multi-modal tracking performance. Our visual adapter effectively integrated auxiliary modalities into the RGB modality by learning robust and adaptable prompts that capture frequency characteristics. Additionally, our memory adapter was designed to store and retrieve global temporal context, ensuring that the most reliable temporal cues are propagated through sequential frames. Extensive experiments demonstrate that our tracker achieves state-of-the-art performance in all multi-modal tracking benchmarks.

### v2

@@ 4.3 Comparison with the State-of-the-Art Methods
To evaluate the impact of different components on the tunable parameter and model complexity, we conduct a comparison of tunable parameters on the LasHeR dataset. As shown in Table 6, the first row represents the baseline model, where all components except the patch-embedding layer are frozen, resulting in a total of 93.4M parameters, with 1.9M being tunable. The second row shows the result of making all components of the baseline tunable, which improves the tracking performance by 4.3% in precision rate and 3.7% in success rates; however, the number of tunable parameters rises to 93.4M. In the third row, the proposed visual adapter is added to the baseline and made learnable, adding only 1.9M tunable parameters, while improving the tracking performance by 5.9% in precision rate and 4.7% in success rate, compared to the baseline. The last row shows the final version of the proposed method, which has 98.9M model parameters and 7.3M of them are tunable. In addition, our method introduces about 1 GFlops computational cost more than the baseline.
4.8 Comparison of Visual Adapters.
@@
We conduct comparison experiments between our proposed visual adapter and the widely adopted Fovea fusion adapter [24, 47, 59], fine-tuning both under identical settings. The “–” row in Table 5 represents a strong baseline without any dedicated adapter. Introducing the Fovea fusion adapter improves tracking performance on the LasHeR dataset, with precision and success rates increasing from 65.9% and 51.8% to 69.0% and 54.4%, respectively. Similar trends are observed on DepthTrack and VisEvent, where Fovea yields modest but consistent gains, indicating that spatial re-weighting of features contributes positively to performance. Replacing Fovea with our frequency-aware visual adapter leads to further improvements, achieving 71.8% precision and 56.5% success on LasHeR. This represents absolute gains of +5.9% (PR) and +4.7% (SR) over the baseline, and +2.8% / +2.1% over Fovea. A similar trend is seen on DepthTrack, where the F-score improves from 59.8% (no adapter) and 60.4% (Fovea) to 61.3%. On VisEvent, precision rises from 78.4% (baseline) and 79.0% (Fovea) to 79.5%. Although our module introduces 3.8 M tunable parameters in total, this is only 1.7 M more than the Fovea variant and remains lightweight compared with full fine-tuning of the backbone. The extra capacity is used to model frequency, domain interactions across RGB and auxiliary modalities, which explains why the gain is most pronounced on LasHeR, where motion blur and illumination variation make high-frequency cues especially informative, while still yielding steady improvements on the other two datasets. In sum, Table 5 confirms that our adapter offers the best accuracy–efficiency trade-off among the tested designs, outperforming both the plain baseline and the strong Fovea fusion across all metrics without incurring a prohibitive parameter cost.
@@
4.9 Qualitative Comparison and Feature Analysis.
@@
2) Feature Representation Analysis: To provide a more comprehensive visual analysis, we further inspect the feature visualization produced by both the visual and memory adapters. For the visual adapter, Figure 8 arranges three columns: the left column shows the raw search region, the middle column depicts the feature map generated by the baseline without our adapter, and the right column presents the features after using the adapter. Once the adapter is enabled, the target contour becomes crisp and the clutter in the background is markedly suppressed, confirming that frequency-aware recalibration strengthens discriminative cues while filtering out irrelevant noise. The memory adapter cannot be plotted directly because it stores one-dimensional vectors, so we instead compare the target‐region feature maps before and after the memory adapter in Figure 9. After memory aggregation, the foreground activations are noticeably sharper and the separation from the background is clearer, illustrating that temporal information preserved in memory further refines spatial representations and stabilizes tracking.
5 Conclusion
In this paper, we proposed a novel prompt-learning framework that learns frequency- and memory-aware prompts for multi-modal object tracking. The frequency-guided visual adapter adaptively fuses auxiliary modalities with RGB by exploiting complementary spatial, channel, and frequency information, while the multi-level memory adapter stores and retrieves temporal cues to ensure consistent long-range propagation across frames. This dual-adapter design enables efficient and robust prompt learning on frozen trackers. Extensive experiments on RGB-T, RGB-D, and RGB-E benchmarks demonstrate that our method achieves state-of-the-art performance, significantly outperforming both fully fine-tuned and adapter-based baselines.
Acknowledgment
This work was supported by the National Natural Science Foundation of China (62072232, 62576098), the Key R&D Project of Jiangsu Province (BE2022138), the Fundamental Research Funds for the Central Universities (021714380026), the program B for Outstanding Ph.D. candidates of Nanjing University, and the Collaborative Innovation Center of Novel Software Technology and Industrialization. (Boyue Xu and Ruichao Hou contributed equally to this work.)